OpenStudy (anonymous):

Derivitive for e^(-2t)?

OpenStudy (anonymous):

do you know chain rule?

OpenStudy (kropot72):

\[-2e ^{-2t}\]

OpenStudy (anonymous):

yeah i do. i think the answer is -2e^(-2t)

OpenStudy (anonymous):

but its not working in the problem i have so think i'm doing something wrong here.

OpenStudy (anonymous):

\[\LARGE (e^u)'=e^u\cdot u'\] leads you in the answer you got. or.. \[\LARGE -\frac{2}{e^{2t}}\]

OpenStudy (anonymous):

(fog)' = (f'og) x g'

OpenStudy (anonymous):

ok thank you very much :)

OpenStudy (anonymous):

you're welcome...